mathslover (mathslover):

hello every one . I am posting a tutorial on : The topic of quadratic equation : roots of quadratic equation

mathslover (mathslover):

So , Today i am going to post a tutorial on A theorem on quadratic equations : Have you ever thought of that is their any proof of : Can a quadratic equation only have 2 roots , cant it more ? There is a simple reason that is the degree of polynomial is 2 . But I am going to post the exact proof for this : (* 2 distinct real roots ) Quadratic equation general formula ----------> \(\large{ax^2+bx+c=0}\) So let us have : three roots : \(\alpha\) , \(\beta\) , \(\gamma\) : alpha , beta and gamma are distinct roots and none of them equals zero . If alpha , beta and gamma are the roots then it is necessary that: \(\large{a(\alpha^2)+b(\alpha)+c=0}\) ------------(1) \(\large{a(\beta^2)+b(\beta)+c=0}\) ------------(2) \(\large{a(\gamma^2)+b(\gamma)+c=0}\) ------------(3) So first of all we will subtract eq. 2 from eq. 1 that is : \[\large{a(\alpha^2)+b(\alpha)+c-a(\beta^2)-b(\beta)-c=0}\] \[\large{a(\alpha^2)-a(\beta^2)+b(\alpha)-b(\beta)+c-c=0}\] \[\large{a(\alpha^2-\beta^2)+b(\alpha-\beta)=0}\] \[\large{a(\alpha+\beta)(\alpha-\beta)+b(\alpha-\beta)=0}\] Applying inverse of distributive property we get : \[\large{(\alpha-\beta)[a(\alpha+\beta)+b]=0}\] \(\large{a(\alpha+\beta)+b=0}\)---------------(4) SINCE : we have taken alpha and beta as two distinct roots , hence their diff. can not be 0 . Now similarly subtracting eq. (3) from eq. (2) \[\large{a(\gamma^2)+b(\gamma)+c-a(\beta^2)-b(\beta)-c=0}\] \[\large{a(\gamma^2)-a(\beta^2)+b(\gamma)-b(\beta)+c-c=0}\] \[\large{a(\gamma^2-\beta^2)+b(\gamma-\beta)=0}\] \[\large{a(\gamma+\beta)(\gamma-\beta)+b(\gamma-\beta)=0}\] \[\large{(\gamma-\beta)[a(\gamma+\beta)+b] = 0}\] \(\large{a(\gamma+\beta)+b=0}\) ------------(5) SINCE : We have taken gamma and beta as two distinct roots , hence their diff. can not be 0 Now Subtracting eq.(5) from eq.(4) : \[\large{a(\alpha+\beta)+b-a(\gamma+\beta)-b=0}\] \[\large{a(\alpha+\beta)-a(\gamma+\beta)+b-b=0}\] \[\large{a(\alpha+\beta-\gamma-\beta)=0}\] \[\large{a(\alpha-\gamma)=0}\] But this is not possible : Because alpha and gamma are distinct roots and alpha is not equal to zero . So , their product can not be zero . Thus the assumption that a quadratic equation has three distinct real roots is wrong . Hence a quadratic equation can not have more than 2 roots.

OpenStudy (anonymous):

awesome @mathslover :)

OpenStudy (unklerhaukus):

and \(a\) cannot be zero or else the equation would not be a quadratic

mathslover (mathslover):

yes thanks for that point also @UnkleRhaukus

OpenStudy (unklerhaukus):

Great work mathslover this tutorial is goood/.

mathslover (mathslover):

thanks @annas and @UnkleRhaukus I will post some more tutorials soon whenever i will get time

OpenStudy (anonymous):

welcome maths :) wish you best of luck for future :)

OpenStudy (anonymous):

How do you guys post such long tutorials so quickly?

mathslover (mathslover):

i just make it ready and then post this

OpenStudy (anonymous):

I see ... that's actually quite a good idea.

mathslover (mathslover):

Good to hear that u r also making some tutorials best of luck : thanks also

mathslover (mathslover):

So friends here is my whole tutorial in a drawing form : i did it on my own

mathslover (mathslover):

@dpaInc , @eashmore , @jhonyy9 have a look to this tutorial :)

mathslover (mathslover):

@vishweshshrimali5 and @Zarkon sir please give some suggestions

OpenStudy (preetha):

That is impressive.

mathslover (mathslover):

Thanks mam .. I will post some more soon .. i just need support from u all :)

OpenStudy (preetha):

KV rocks!

OpenStudy (preetha):

Kendriya Vidyalaya

mathslover (mathslover):

Thanks a lot mam ! .. Thanks to my teachers and to all of u also .... :)

OpenStudy (anonymous):

Your equation \(a(\alpha-\gamma)=0\) simply implies \(\alpha=\gamma\), thus confirming there are only two roots: \(\alpha\) and \(\beta\). (This is just how I looked at it. After reading your entire tutorial, your method makes sense.)

mathslover (mathslover):

Yes @Limitless thanks a lot for adding this to my tutorial ... this thing looks good : \[\huge{\alpha-\gamma=0}\] \[\huge{\alpha=\gamma}\] Only if alpha is not equal to 0 that is given to us.. thanks again

OpenStudy (anonymous):

It doesn't matter what \(\alpha\) is; it's true that \(\alpha=\gamma\) regardless.

mathslover (mathslover):

K ! thanks

OpenStudy (anonymous):

Really nice work.............Well done!

mathslover (mathslover):

thanks a lot @nitz

OpenStudy (anonymous):

i must say your tutorials are really nice and useful!

mathslover (mathslover):

http://openstudy.com/study#/updates/4ff50f22e4b01c7be8c837ca have a look to my new tutorial

OpenStudy (goformit100):

Nice

mathslover (mathslover):

:p medals level :)